### v2

# Dolphin v1.0 Technical Report

Abstract
Ultrasound is vital in modern medicine but faces challenges like operator dependence, image noise, and dynamic real-time scanning, which hinder effective AI integration. Though large multimodal models excel in other medical imaging areas, they struggle with ultrasound’s unique complexities. To bridge this gap, we introduce Dolphin v1.0 (V1) and its reasoning-augmented version, Dolphin R1—the first large-scale multimodal ultrasound foundation models designed to unify diverse clinical tasks within a single vision-language framework. To tackle the variability, noise, and operator dependence inherent in ultrasound imaging, we curated a groundbreaking multimodal dataset at the 2-million-scale, integrating textbook knowledge, public ultrasound data, synthetic knowledge-distilled samples, and general multimodal corpora. This comprehensive dataset ensures the models achieve robust perception, strong generalization, and broad clinical adaptability across medical domains.
The Dolphin series adopts a progressive three-stage training strategy that combines domain-specialized pretraining, instruction-driven alignment, and reinforcement-based refinement. The baseline Dolphin v1.0 establishes reliable performance in classification, detection, regression, and report generation, offering a strong general-purpose model for clinical ultrasound analysis. Building upon this foundation, Dolphin R1 leverages reinforcement learning with verifiable ultrasound-specific reward signals, thereby enabling deeper diagnostic inference, enhanced reasoning transparency, and more interpretable decision pathways.
To systematically evaluate performance,We adopt U2-Bench for evaluation eight representative ultrasound tasks, ranging from lesion localization and organ detection to clinical value estimation and structured report generation. Experimental results demonstrate that Dolphin R1 achieves a U2-score of 0.5835—more than twice the score of the second-best model (0.2968)—establishing a new state of the art in multimodal ultrasound understanding. Meanwhile, Dolphin v1.0 also delivers competitive performance, validating the effectiveness of the unified training framework. Further comparison between standard mode (V1) and deep reasoning mode (R1) reveals that reasoning-enhanced training significantly boosts diagnostic accuracy, consistency, and interpretability. These findings underscore the importance of integrating reasoning into foundation models for high-stakes medical domains, providing a scalable paradigm for future developments in intelligent ultrasound analysis.
Introduction
Ultrasound (US) is one of the most widely utilized imaging modalities in clinical practice, playing a critical role in obstetrics, cardiology, emergency medicine, and resource-limited settings due to its real-time imaging capability, portability, and cost-effectiveness(Hewson & Bedforth (2023)). However, despite its broad applicability, ultrasound interpretation remains highly challenging compared to other medical imaging modalities such as computed tomography (CT), magnetic resonance imaging (MRI), and positron emission tomography (PET). These modalities provide high spatial resolution, standardized anatomical views, and consistent image quality, whereas ultrasound is inherently operator-dependent, susceptible to imaging artifacts, and exhibits significant variability in acquisition quality(Sharma et al. (2021)).
Currently, multimodal large models have emerged as a powerful paradigm in artificial intelligence research(OpenAI (2023), Bai et al. (2025), Lin et al. (2023), Wu et al. (2024)), demonstrating remarkable capabilities in processing and integrating information across diverse modalities such as text, images, audio, and video. Currently, Multimodal large models have demonstrated significant potential in the medical field, offering advanced capabilities for integrating and analyzing diverse data types, including medical images (MRI,CT,PET; Lee et al. (2024); Volkov et al. (2025); Shui et al. (2025)), electronic health records(AlSaad et al. (2024)), genomic data(Fallahpour et al. (2025)), and clinical notes(Han et al. (2024)). While these models have achieved remarkable success in processing static, high-resolution medical images and structured clinical data, their application to ultrasound imaging presents unique challenges that remain largely unaddressed.
The translation of multimodal approaches to ultrasound is hindered by several critical limitations. First, unlike the standardized acquisition protocols for CT or MRI, ultrasound examinations are highly operator-dependent, resulting in significant variability in image quality and anatomical views (Stasi & Ruoti (2015)). Second, the dynamic nature of ultrasound video sequences - often containing crucial temporal information about tissue motion and blood flow - is fundamentally different from the static images that current multimodal models typically process (Szabo (2013)). Third, the characteristic speckle noise and lower signal-to-noise ratio of ultrasound images compared to other modalities create additional challenges for feature extraction and interpretation (Loizou & Pattichis (2022)). Furthermore, existing multimodal models struggle with the sparse and often subjective nature of ultrasound reports, which typically lack the detailed anatomical descriptions found in radiology reports for other imaging modalities(Wang et al. (2024)). Given the aforementioned challenges, developing large multimodal models specifically tailored for ultrasound imaging has become an urgent necessity. In recent years, several studies have attempted to apply ultrasound images in the research of large-scale multimodal models. The EchoCLIP model proposed by Christensen et al. (2024) learns the correspondence between cardiac ultrasound images and expert-annotated cardiac case reports, thereby achieving an understanding of cardiac-related semantic information. Subsequently, Guo et al. (2024) proposed LLaVA-Ultra, which integrates the CLIP and SAM encoders and constructs a general-purpose large model for the ultrasound domain through domain-specific fine-tuning. It has outperformed LLaVA-Med(Li et al. (2023)) on benchmark tests across various general medical tasks.
Although multimodal large models have been increasingly explored in the ultrasound domain, existing approaches often lack comprehensive generalization and in-depth medical reasoning capabilities. To address these challenges, we introduce Dolphin V1 and Dolphin R1, the world’s first general-purpose multimodal ultrasound large models. Dolphin V1 is a vision-language model capable of jointly understanding ultrasound images, videos, dynamic scanning sequences, and associated clinical text, while Dolphin R1 further enhances the model’s reasoning and diagnostic inference capabilities.
To support model development, we curated a large-scale multimodal ultrasound dataset containing over 2 million high-quality samples sourced from medical textbooks, public ultrasound repositories, general multimodal data, and knowledge-distilled synthetic data. This diverse and extensive dataset underpins the models’ robustness and clinical adaptability. Additionally, we propose the Dolphin Ultrasound Data Protocol (DUDP), a standardized data interface that unifies and formalizes a wide range of ultrasound tasks, laying the groundwork for future research and development in this domain.
In our work, we first unify a wide range of complex ultrasound tasks by abstracting them into four core task categories: classification, detection, regression, and generation. To effectively train the model, we employ a progressive three-stage training strategy, which simultaneously emphasizes the model’s generalization capability and medical understanding. Furthermore, we incorporate large-scale reinforcement learning based on ultrasound data, endowing the model with advanced reasoning abilities for complex diagnostic scenarios.
The experimental results indicate that our model outperforms all compared methods across the eight diverse ultrasound tasks defined in the U2-Bench benchmark. It achieves superior performance on a wide range of tasks, including disease diagnosis, organ detection, and medical report generation, demonstrating robustness and generalization capability in both low-level perception and high-level reasoning. Notably, our model achieves a U2-Bench score of 0.5835, significantly surpassing the second-best performing model, Gemini-2.5 Pro, which scores 0.2968. This result establishes a new state of the art in ultrasound-based multimodal modeling.
Our key contributions can be summarized as follows:
  • We construct a large-scale, multimodal ultrasound dataset comprising over 2 million high-quality samples, covering both supervised fine-tuning (SFT) and reinforcement learning (RL) data.
  • We propose a progressive three-stage training framework that effectively enhances the model’s generalization capability and medical reasoning.
  • Our model achieves state-of-the-art performance on the U2-Bench benchmark, significantly outperforming existing methods across a wide range of ultrasound tasks, demonstrating strong low-level perception and high-level reasoning abilities.
These advancements collectively pave the way for significant progress in ultrasound-based medical AI, fostering more accurate, efficient, and intelligent clinical decision-making.
Data Curation
2.1 Post-training data
To enable the model to possess expert-level ultrasound image recognition capabilities, we constructed a large-scale, high-quality ultrasound image dataset. Specifically, our data is primarily sourced from the following four aspects: (1) Ultrasound textbooks and clinical guidelines - to enhance the model’s fundamental ultrasound recognition capabilities; (2) Publicly available ultrasound datasets covering classification, segmentation, detection, measurement, and other tasks - to improve the model’s disease diagnostic abilities; (3) General medical data - to enhance the model’s generalization and conversational capabilities; (4) Distilled knowledge data - to strengthen the model’s instruction-following and formatted output abilities.
During this process, we also constructed synthetic ultrasound data through the following approaches: (1) Question templates – Generated deterministic responses for structured data construction. (2) VQA data – Created using GPT-generated plausible options for ultrasound visual question answering. (3) Knowledge distillation data – Derived ultrasound datasets through model-based knowledge transfer, following the methodology of Xu et al. Xu et al. (2023a) on Evol-Instruct to iteratively enhance the quality of distilled data.
Our data generation pipeline is meticulously designed to minimize the risk of hallucination and ensure high fidelity in clinical representation. The process begins with a well-defined task framework and template construction, which is followed by an initial review by board-certified physicians to validate the clinical relevance and diagnostic coherence of the proposed structure. Upon achieving expert consensus, synthetic data is generated in accordance with the predefined templates. This synthesized data then undergoes a second round of rigorous evaluation by medical professionals, who provide detailed feedback to guide iterative refinement and correction. Only after substantial alignment with clinical standards is achieved does the data proceed to a final validation stage, where it is independently assessed by physicians for inclusion in the training corpus. In total, this systematic approach incorporates three iterative rounds of cross-validation, ensuring robust clinical plausibility, diagnostic accuracy, and methodological rigor throughout the data creation process.
Figure 1: Distribution of the Dolphin Ultrasound Large-Scale Dataset
2.1.1 Ultrasound teaching material data
We collected a large number of classic ultrasound textbooks and industry guidelines, uniformly converted these materials into Markdown format, manually filtered out non-substantive content such as covers, back covers, and margins, and corrected grammatical errors and formulas that were inaccurately converted. Based on these ultrasound textbooks, we constructed the following three datasets: the Ultrasound Question-Answering Dataset, the Ultrasound Caption Dataset, and the Ultrasound Image-Text Explanation Dataset.
Ultrasound Question-Answering Dataset: We segmented the textbook data into chunks and employed GPT to generate contextually appropriate questions for each segment. Through self-distillation, we regenerated corresponding answers to ensure strict question-answer consistency.
Ultrasound Caption Dataset: The ultrasound textbook corpus contains systematic image-caption pairs, which we processed into a structured dataset through automated format matching to ensure precise image-annotation correspondence. It should be noted that our constructed caption dataset contains cases with multiple images, which helps enhance the model’s multi-image comprehension capability.
Ultrasound Image-Text Explanation Dataset: The ultrasound textbooks contain detailed textual descriptions of medical images. We first identify the text passages corresponding to each image, then process these descriptions through GPT using self-distillation to generate high-quality question-answer pairs. This approach ensures data accuracy while simultaneously enhancing the model’s response diversity and instruction-following capabilities.
2.1.2 Public ultrasound datasets
To enhance the model’s diagnostic capabilities across various tasks, we collected a substantial number of publicly available ultrasound datasets and established a unified data processing pipeline. We developed a unified file standard called Ultrasound Data Protocol (UDP) to harmonize diverse ultrasound tasks by converting them into a standardized JSON format. The protocol comprehensively handles diverse ultrasound tasks, supporting multimodal data types including ultrasound images, video sequences, segmentation masks, and associated quantitative measurements.
Specifically, we categorize ultrasound tasks into the following standardized classes: (1) Anatomical Recognition (2) Standard View Identification (3) Diagnostic Classification (4) Tissue Segmentation (5) Biometric Measurement. Corresponding ultrasound question-answering datasets were constructed for each task category. To mitigate hallucination risks, we developed over a hundred fixed question templates for each task category. For each sample, we randomly selected a question template and populated it with JSON-formatted content, thereby constructing a high-quality ultrasound question-answering dataset. Figure 2 presents an authentic diagnostic data sample from our constructed dataset.
Furthermore, to ensure data diversity, we employed GPT for partial data augmentation and utilized another LLM to evaluate response quality, guaranteeing that the JSON-formatted content remains embedded within the generated responses.
Figure 2: Overview of the Dolphin Ultrasound Public Data Pipeline. The framework includes data collection, expert-guided filtering, and standardization via the Dolphin Ultrasound Data Protocol (DUDP), ensuring consistency across multimodal sources. Tasks are unified into four categories—classification, detection, regression, and generation—covering diverse clinical scenarios. The dataset distribution demonstrates balanced coverage across major anatomical regions, supporting robust model generalization.
2.1.3 General data
To enhance the model’s general capabilities while incorporating ultrasound-specific knowledge, we selected corresponding public datasets based on different competency requirements. The selected public datasets are primarily utilized to enhance the following model capabilities: (1) Comprehension; (2) Conversational; (3) Deep reasoning; (4) Medical; (5) Image-text dialogue; The primary public datasets adopted in our study are as follows: LLaVA-One-Vision. (Li et al., 2024), AutoIF. (Dong et al., 2024), Alpaca-GPT4 (Liu et al., 2023), Medical-O1-Reasoning (Chen et al., 2024). These datasets have been meticulously curated to comprehensively enhance the model’s performance across multiple dimensions.
2.1.4 Distilled knowledge data
Following the methodology of (Xu et al., 2023b), we first constructed a set of ultrasound data question templates. The difficulty level of questions was enhanced through depth-wise augmentation, while their diversity was improved via breadth-wise expansion. All question categories were strictly standardized within the ultrasound medical domain, and we distilled ultrasound-specific data from both GPT and DeepSeek. Additionally, we designed a data screening pipeline that employs large language models to assess the logical consistency of outputs, ultimately yielding a high-quality distilled dataset.
Figure 3: Examples of constructed ultrasound question–answer pairs. The dataset integrates diverse clinical tasks, including fetal abdominal view assessment, lung ultrasound severity scoring, and soft tissue lesion evaluation, with corresponding expert-verified answers. These QA pairs provide structured supervision signals that enhance both low-level image understanding and high-level diagnostic reasoning in Dolphin models.
2.2 Data Filtering
To ensure the quality and reliability of the distilled dataset, we implemented a rigorous data filtering pipeline consisting of multiple stages:
  • Model-Based Semantic Filtering: Employed LLM-based scoring (e.g., GPT-4) to assess:
    • Relevance: Whether answers properly address the questions.
    • Accuracy: Factual correctness based on medical knowledge.
    • Consistency: Logical coherence within the response.
  • Rule-Based Filtering:
    • Content consistency: Filtering incorrectly generated samples through character matching in JSON files or textbook captions.
    • Format screening: Filter out samples with overly short responses or excessive numbers of images, while also removing sensitive information
  • Expert Validation: Assess the rationality of dataset construction through medical expert evaluation, and eliminate unreasonable datasets.
Through the aforementioned data filtering process, we obtained a high-quality medical ultrasound dataset that covers textbooks and publicly available data, encompassing multiple anatomical regions and diverse diagnostic tasks.
Model Training
Dolphin v1.0 is built upon the Qwen2.5-VL model architecture, which represents an advanced vision-language foundation model developed by (Bai et al., 2025). The system enhances the original framework with a redesigned Vision Transformer (ViT) for optimized visual processing.
We systematically trained two model variants: a base version with 7 billion parameters and an extended version with 72 billion parameters. For each scale, we optimized the training configurations separately, with particular emphasis on investigating the impact of training epochs on both the generalizability and domain-specific performance of the models.
3.1 Training Recipe
The Dolphin v1.0 training pipeline follows a three-stage progressive framework, designed to integrate domain-specific knowledge, align with human preferences, and refine autonomous decision-making capabilities. The specific three stages are as follows: (1) Domain-Specialized Training; (2) Instruction-Driven Alignment; (3) Autonomous Reinforcement Refinement. Through two key phases - domain knowledge injection and instruction alignment - we developed the Dolphin V1.0 model. Building upon this foundation, we conducted reinforcement learning training using extensive ultrasound image-text data to obtain the Dolphin R1 model. Our subsequent research further explored the significance and impact of reinforcement learning on large-scale medical models.
Figure 4: Training pipeline of the Dolphin series model. The process consists of three key stages: domain-specific pretraining on 2M multimodal medical data, instruction finetuning with 10.3k curated samples to enhance alignment with clinical tasks, and self-play reinforcement learning on 60.7k interactions to further improve reasoning and decision-making ability.
3.1.1 Domain-Specialized Training
The goal of the Domain-Specialized Training is to inject ultrasound-specific knowledge into the large language model while preserving its inherent generalizability to the greatest extent possible. During this phase, the model is trained on extensive textbook-based ultrasound data and publicly available ultrasound datasets. The training encompasses ultrasound images of 15 major anatomical systems, with particular emphasis on developing fundamental capabilities in disease diagnosis, anatomical localization, and scan plane recognition. Notably, we transform lesion localization information into standardized directional descriptors (e.g., ”upper left quadrant”), thereby endowing the model with precise pathological assessment capabilities.
Recent studies conducted by (Chen et al., 2020), (Yang et al., 2025) have robustly evidenced that domain-specific fine-tuning of large language models with medical datasets may substantially compromise their core linguistic competencies, resulting in pronounced catastrophic forgetting phenomena. This empirical finding underscores the critical need for balanced adaptation strategies when specializing foundational models for medical applications.
During our model training process, we encountered the same issue. Therefore, when fine-tuning the model with ultrasound data using SFT (Supervised Fine-Tuning), we incorporated data from other domains, with particular emphasis on the model’s fundamental capabilities in the following aspects: reasoning, multi-turn dialogue, instruction-following , OCR detection, and mathematical reasoning. The introduction of general-domain data helped maintain these core competencies.
Interestingly, we observed that as the volume of both ultrasound-specific data and general-domain data increased proportionally, the model demonstrated simultaneous improvements in both ultrasound-related tasks and general-domain performance. This synergistic enhancement suggests that domain-specific specialization and general capability maintenance can be mutually reinforcing when balanced appropriately during training.
3.1.2 Instruction-Driven Alignment
During the Instruction-Driven Alignment phase, our primary focus is refining the model’s output to ensure strict adherence to predefined formats and content requirements. This involves fine-tuning the model’s responses to align with structured templates, domain-specific conventions, or user-specified guidelines. Our methodology combines automated techniques with expert human oversight:
  • Distilled Knowledge Integration:
    • We constructed a small-scale instruction dataset primarily derived from distilled knowledge data. This dataset captures high-quality, structured responses from existing aligned models. It enables the model to learn and replicate the desired stylistic patterns, ensuring consistency with established outputs while minimizing bias.
  • Expert-Guided Refinement:
    • Submit first-phase model outputs to domain specialists (e.g., medical professionals) for review.
    • Incorporate expert feedback through Direct annotation of inaccuracies.
    • Implement iterative refinement cycles to progressively improve output quality
At this stage, to prevent degradation of the model’s inherent capabilities, we adopted a relatively small learning rate and performed only 1 epoch of fine-tuning on the constructed Instruction-Driven Alignment dataset.
3.1.3 Autonomous Reinforcement Refinement.
Recent breakthroughs in reasoning models, such as OpenAI’s o-series ((Jaech et al., 2024)) and DeepSeek-R1 ((Guo et al., 2025)), have demonstrated remarkable performance in complex tasks, largely due to innovations in reinforcement learning (RL) techniques. Among these, Group Relative Policy Optimization (GRPO; (Shao et al., 2024)) has gained prominence for its efficiency in optimizing model behavior through verifiable reward signals. Unlike traditional supervised fine-tuning (SFT), which often suffers from answer overfitting and shortcut learning(Hao et al., 2025), GRPO encourages models to explore reasoning pathways autonomously, improving robustness and generalization—particularly critical in high-stakes domains like medicine(Pan et al., 2025).
For GRPO model training, verifiable results are of paramount importance. Each sample within our meticulously constructed UDP dataset is annotated with precise categories and measurement outcomes to guarantee data reliability and reproducibility. For each ultrasound task, we systematically generate corresponding questions based on the desired output format. The constructed dataset is categorized into two primary types:
  • General Responses: Open-ended answers requiring descriptive or analytical output.
  • Multiple-Choice Responses (A/B/C/D): Structured selections with rigorously controlled proportions to ensure balanced representation.
During the training process, we adopted the configuration proposed by (Guo et al., 2025), setting the reward function to consist of format reward and outcome reward, both assigned a value of 1.0. Through our exploration, we found that the model’s performance during the RL phase largely depends on the fundamental capabilities acquired in the SFT phase. The GRPO training approach primarily serves to activate the model’s inherent capabilities rather than inject new knowledge into it.
Experiment
To evaluate the ultrasound understanding capabilities of large vision-language models (LVLMs), we adopt U2-Bench (Le et al., 2025), a comprehensive benchmark specifical designed for ultrasound-related tasks. The experimental results presented in Table 1 provide a comprehensive evaluation of various large vision-language models (LVLMs) on the U2-Bench, a benchmark specifically designed to assess ultrasound-related multimodal understanding. Among the models evaluated, the Dolphin series—particularly Dolphin v1.0, Dolphin R1, and their smaller variant Dolphin v1.0-7B—demonstrate competitive performance across multiple tasks, with Dolphin R1 achieving the highest U2-Score of 0.5835 among all models tested.
[…]
Models $D$ VRA $L$ OD KD CVE RG CG U2-Score
Acc. F1 Acc. F1 Acc. Acc. Acc. RMSE MAE %tol BLEU% Rouge% BERT% BLEU% Rouge% BERT%
Random Guessing 0.4143 0.4135 0.3195 0.3184 0.1118 0.0680 0.1120 0.5472 0.4352 18.776 - - - - - - 0.2125
Medical-Specific Open-Source Models
[…]
MedDr 0.4508 0.3118 0.2071 0.1214 0.0720 0.0881 0.0900 0.2144 0.1786 38.2642 2.7998 13.5060 72.2050 33.4939 49.6236 81.2078 0.2373
Open-Source Multimodal Models
Qwen-2.5-VL-3B-Instruct 0.4503 0.3591 0.2097 0.1492 0.0696 0.0649 0.0894 0.5008 0.4519 18.9055 3.5018 15.0327 72.8419 27.6748 44.7618 79.8849 0.2095
Qwen-2.5-VL-7B-Instruct 0.4821 0.3860 0.2181 0.1665 0.0750 0.0704 0.1000 0.4646 0.4337 19.7115 3.7100 15.5600 73.1500 29.4400 47.0000 81.1500 0.2235
Qwen-2.5-VL-3B-Instruct 0.4503 0.3591 0.2097 0.1492 0.0696 0.0649 0.0894 0.5008 0.4519 18.9055 3.5018 15.0327 72.8419 27.6748 44.7618 79.8849 0.2095
Qwen-2.5-VL-7B-Instruct 0.4821 0.3860 0.2181 0.1665 0.0750 0.0704 0.1000 0.4646 0.4337 19.7115 3.7100 15.5600 73.1500 29.4400 47.0000 81.1500 0.2235
Qwen-2.5-VL-32B-Instruct 0.4812 0.3860 0.2864 0.2071 0.1700 0.0755 0.0880 0.3414 0.3015 27.4038 1.1900 13.0100 68.1400 14.7700 38.6800 77.3900 0.2449
Qwen-2.5-VL-72B-Instruct 0.4895 0.4556 0.2559 0.1789 0.1150 0.0660 0.0860 0.3224 0.2733 37.9370 3.0900 15.0600 72.6600 28.1600 44.2800 80.9100 0.2421
[…]
InternVL3-9B-Instruct 0.4447 0.3716 0.1926 0.1083 0.3000 0.1416 0.0940 0.2429 0.1733 50.8738 2.1600 14.7000 72.2100 21.5900 43.1300 80.9800 0.2566
[…]
Phi-4-Multimodal-Instruct 0.3686 0.1148 0.2452 0.0537 0.0350 0.0815 0.1600 0.2249 0.2006 16.1972 3.2700 16.5800 73.2700 3.8700 22.9800 73.0800 0.2168
[…]
Closed-Source Multimodal Models
Doubao-1.5-Vision-Pro-32k 0.5580 0.2597 0.2922 0.2147 0.1700 0.0729 0.1240 0.3664 0.3377 33.1731 0.7100 6.6450 72.4000 8.6400 33.3000 78.4200 0.2587
GPT-4o-Mini 0.4924 0.3784 0.1922 0.1272 0.1357 0.0846 0.0960 0.2267 0.1976 19.2308 4.9400 17.5200 74.1300 11.7300 36.2900 77.5300 0.2388
GPT-4o 0.4928 0.4132 0.1504 0.0974 0.1161 0.0850 0.0840 0.3712 0.3527 15.7895 2.6800 14.7700 73.3500 33.7700 49.9600 81.5800 0.2253
Gemini-1.5-Pro 0.3781 0.2247 0.0909 0.0476 0.2700 0.0661 0.0980 0.2772 0.2205 40.7051 0.5800 9.9400 70.5500 28.5800 45.9200 80.0200 0.1999
[…]
Claude-3.7-Sonnet 0.2121 0.0449 0.1453 0.0479 0.1356 0.0540 0.0760 0.1764 0.1500 36.0215 0.6900 12.2300 68.7400 1.2900 16.6600 71.6600 0.1596
[…]
4.1 Performance Overview
Dolphin R1 excels in classification and detection tasks, particularly in View Recognition and Assessment (VRA), where it achieves the highest accuracy (0.6943) and F1 score (0.5821), indicating strong spatial understanding and anatomical structure recognition capabilities. Similarly, it outperforms other models in Lesion Localization (LL), Organ Detection (OD), and Keypoint Detection (KD), with accuracy scores of 0.4775, 0.6003, and 0.5080 respectively, suggesting robust spatial reasoning and fine-grained visual recognition. These results highlight the model’s ability to handle complex ultrasound-specific visual patterns and anatomical variations. Despite strong performance in classification and detection tasks, Dolphin R1 exhibits limitations in Clinical Value Estimation (CVE) and Report Generation (RG), achieving only moderate accuracy in numerical parameter.
In comparison, Dolphin v1.0 shows slightly lower but still strong performance across these tasks, while Dolphin v1.0-7B, as a smaller variant, exhibits a noticeable performance drop, particularly in OD and KD, indicating the importance of model scale in capturing fine-grained visual features.
4.2 Impact of Model Scale
The comparison between Dolphin v1.0-7B and its larger 72B counterparts—Dolphin v1.0 and Dolphin R1—highlights the significant impact of model scale on multimodal understanding in ultrasound-related tasks. The 72B models consistently outperform the smaller 7B variant across classification, detection, and generation tasks, with the most notable improvements seen in Organ Detection (OD), Keypoint Detection (KD), and Caption Generation (CG), indicating that larger models are better suited to capture the visual complexity and semantic nuances present in medical ultrasound imaging.
Moreover, Dolphin R1 achieves a significantly improved U2-Score compared to Dolphin v1.0, demonstrating that advanced training strategies such as GRPO can effectively enhance model performance. This highlights the importance of both model scale and training methodology in pushing the boundaries of ultrasound image understanding.
4.3 Comparison Between Standard Mode and Deep Reasoning Mode
Figure 5: Comparison of Standard and Deep Reasoning models on DolphinV1.0 classification tasks, evaluated by accuracy, precision, recall, and F1-score.
As shown in Figure 5, we also evaluated the performance of our model, Dolphin R1, on disease diagnosis tasks under both standard mode and deep reasoning mode. The results demonstrate that, in the majority of tasks, enabling the deep reasoning mode leads to a notable improvement in diagnostic accuracy. Moreover, the deep reasoning mode offers a more thorough and detailed analytical process, providing interpretable and transparent reasoning trails. This enhanced interpretability not only supports more informed clinical decision-making but also aligns better with user preferences, as evidenced by the fact that physicians tend to favor the deep reasoning mode in practice.
Conclusion
In this work, we introduced Dolphin v1.0 and Dolphin R1, advancing multimodal ultrasound understanding through large-scale data curation and progressive training strategies. The 200w-scale dataset provided the foundation for robust learning across heterogeneous tasks, while the integration of reinforcement learning in Dolphin R1 endowed the model with superior reasoning and interpretability. Our experimental results show that Dolphin R1 achieved a U2-score of 0.5835, surpassing all competitors by a large margin and doubling the performance of the second-best system, whereas Dolphin v1.0 established a strong baseline in standard mode. The comparative analysis highlights that deep reasoning not only improves quantitative accuracy but also enhances the interpretability of diagnostic processes, aligning more closely with physician preferences. Looking forward, this dual-track design of standard vs. deep reasoning modes provides a flexible paradigm for future medical AI systems, balancing efficiency, accuracy, and clinical usability.
Authors
Core Contributors: Taohan Weng
Contributors [^1] Alphabetical order.: Chi zhang, Chaoran Yan, Siya Liu, Xiaoyang Liu, Yalun Wu, Boyang Wang, Boyan Wang, Jiren Ren, Kaiwen Yan, Jinze Yu, Kaibing Hu, Henan Liu, Haoyun Zheng, Zhenyu Liu, Duo Zhang, Xiaoqing Guo
Project Leader: Hongcheng Guo, Anjie Le
Appendix A Appendix